# GigaSLAM: Large-Scale Monocular SLAM with Hierachical Gaussian Splats

### v1

@@ 4.3 KITTI 360 Dataset
Figure 6: Camera trajectory on the KITTI 360 dataset.Unlike DROID SLAM, which fails due to implicit loop closure, our method explicitly handles large-scale trajectories.
Methods Avg. $0$ 0002 0003 0004 0005 0006 0007 0009 0010
seq. frames 8497 11518 14607 1031 11587 6743 9699 3396 14056 3836
seq. length (m) 6971.088 8403.15 11501.28 1378.73 9975.28 4690.74 7979.92 4887.51 10579.68 3343.49
contains loop -
DROID-SLAM [42] 193.307 110.07 233.87 10.79 169.11 139.02 113.81 577.39 165.34 220.36
Ours (w/o LC) 61.402 25.53 56.59 29.13 116.15 20.51 28.52 203.02 16.83 56.35
Ours (w/ LC) 47.107 17.72 34.98 27.55 40.54 20.46 19.12 197.33 13.22 53.03
Table 3: Camera Tracking Results (ATE RMSE [m] $\downarrow$) on the KITTI 360 Dataset.
@@
To further evaluate the robustness of our method in extensive outdoor scenarios, we conducted an experiment on the KITTI 360 [16] dataset. Unlike the KITTI [10] dataset, where sequence lengths peak at 4,661 frames with an average of approximately 2,109 frames, KITTI 360 sequences are significantly longer, averaging 8,497 frames and reaching a maximum length of 14,607 frames. These extended trajectories introduce unique challenges in monocular RGB-based SLAM, particularly due to the accumulation of errors over such long sequences, which can severely impact tracking accuracy and mapping fidelity.
@@
Notably, almost no existing monocular RGB-based SLAM or VO system has been fully evaluated on KITTI 360 to date. Our method, however, demonstrated the capability to process these ultra-long sequences effectively, providing stable and continuous camera pose estimations across the full length of each sequence. Initial results (Figure 6 and Table 3) indicate acceptable performance on KITTI 360, highlighting our system’s resilience in mitigating error accumulation over extended trajectories. This underscores the scalability and robustness of our approach for unbounded, long-sequence outdoor SLAM tasks.
@@ 5 Conclusion
We present GigaSLAM, the first SLAM system for long-term, kilometer-scale outdoor sequences with monocular RGB input. With a hierarchical sparse voxel structure and metric depth module for precise pose estimation, GigaSLAM offers efficient mapping and robust tracking. Tested on KITTI and KITTI 360, it demonstrates competitive performance and scalability for large-scale SLAM.
@@ A.1 DROID-SLAM based Methods
This small optimization window amplifies the accumulation of scale drift, as errors that could otherwise be corrected through a larger window of jointly optimized frames remain unaddressed. Figure A.1 illustrates this limitation by comparing the co-visibility graphs of DROID-SLAM on KITTI, KITTI 360 and Replica [37]. The co-visibility graph for KITTI reveals sparse connections, primarily linking each frame to its nearest 4–5 neighbors. In contrast, the graph for the indoor Replica dataset is significantly denser, reflecting DROID-SLAM’s inherent suitability for structured, small-scale environments.

### v2

@@ 4.3 KITTI 360, 4 Seasons & A2D2 Dataset
Figure 8: (Left) Camera trajectory on the KITTI 360 dataset. Unlike DROID SLAM, which fails due to implicit loop closure, our method explicitly handles large-scale trajectories. (Right) Our method on A2D2 dataset.
@@
To further evaluate the robustness of our method in extensive outdoor scenarios, we conducted an experiment on the KITTI 360 [18], 4 Seasons[55] & A2D2[12] dataset. Unlike the KITTI [11] dataset, where sequence lengths peak at 4,661 frames with an average of approximately 2,109 frames, KITTI 360 sequences are significantly longer, averaging 8,497 frames and reaching a maximum length of 14,607 frames. So do 4 Seasons and A2D2 datasets that have an average of 20,000 frame input. These extended trajectories introduce unique challenges in monocular RGB-based SLAM, particularly due to the accumulation of errors over such long sequences, which can severely impact tracking accuracy and mapping fidelity.
@@
Notably, almost no existing monocular RGB-based SLAM or VO system has been fully evaluated on these datasets to date. Our method, however, demonstrated the capability to process these ultra-long sequences effectively, providing stable and continuous camera pose estimations across the full length of each sequence. Initial results (Figure 8 and Table 2) indicate acceptable performance on ultra-long sequences, highlighting our system’s resilience in mitigating error accumulation over extended trajectories. On the 4 Seasons dataset, the substantial increase in the number of input frames leads to a dramatic rise in memory requirements, rendering DROID-SLAM inapplicable due to GPU memory exhaustion. As shown, DROID-SLAM runs out of memory on most sequences. Even in the two sequences where it is able to run, our method significantly outperforms it. Although our approach demonstrates strong performance on these sequences, potential loop closures were not detected due to dataset limitations (i.e., DBoW failed to find matches). Nevertheless, our system still maintains reasonable tracking performance under such challenging conditions. This underscores the scalability and robustness of our approach for unbounded, long-sequence outdoor SLAM tasks.
@@ 5 Conclusion
We present GigaSLAM, the first SLAM system for long-term, kilometer-scale outdoor sequences using monocular RGB input. By employing a hierarchical sparse voxel structure and a metric depth module, GigaSLAM enables efficient large-scale mapping and robust pose estimation. Evaluated on the KITTI, KITTI-360, 4 Seasons and A2D2 datasets, our system demonstrates strong performance and good scalability for outdoor SLAM tasks. Looking ahead, future work will focus on improving loop closure detection, particularly under high-speed motion. Enhancing system stability in such scenarios will be key to advancing reliable SLAM for ultra-large-scale, real-world applications. Moreover, extending GigaSLAM to operate under more dynamic conditions and diverse outdoor environments could further improve its practicality.
@@ A.1 DROID-SLAM based Methods
This small optimization window amplifies the accumulation of scale drift, as errors that could otherwise be corrected through a larger window of jointly optimized frames remain unaddressed. Figure A.1 illustrates this limitation by comparing the co-visibility graphs of DROID-SLAM on KITTI, KITTI 360 and Replica [40]. The co-visibility graph for KITTI reveals sparse connections, primarily linking each frame to its nearest 4 to 5 neighbors. In contrast, the graph for the indoor Replica dataset is significantly denser, reflecting DROID-SLAM’s inherent suitability for structured, small-scale environments.